# Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff

### v1

@@ Appendix E Training Details
The jittering noise variance for each method is selected based on performances of both robustness and accuracy. In input jittering training, the three methods in order of 2D denoising, seismic deconvolution and MRI reconstruction use the noise variance of 0.01, 0.05 and 0.05. In SGD jittering training, we choose the DGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
Appendix F Training Speed
We also compare the training speed among different training schemes. All methods are trained on the same device. Training batch sizes for the 2D denoising problem, seismic deconvolution and MRI are 256, 16, and 4 respectively. AT is trained by projected gradient descent. For the 2D denoising problem, $\epsilon=0.01$ with 50 PGD iterations and a stepsize of 0.05. For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 PGD iterations and a stepsize of 0.1.
@@ Appendix F Training Speed
AT is significantly slower than other methods. Notice that SGD jittering is very time efficient, even with the additional noise sampling steps.

### v2

@@ Appendix G Training Details
The jittering noise variance for each method is selected based on performances of both robustness and accuracy. In input jittering training, the three methods in order of 2D denoising, seismic deconvolution and MRI reconstruction use the noise variance of 0.01, 0.05 and 0.05. In SGD jittering training, we choose the SGD jittering variance for each task with 0.01, 0.1 and 0.01 respectively.
Appendix H Training Speed
We also compare the training speed among different training schemes for GD and PGD LU architectures. All methods are trained on the same device. Training batch sizes for the 2D denoising problem, seismic deconvolution and MRI are 256, 16, and 4 respectively. AT is trained by projected gradient descent. For the 2D denoising problem, $\epsilon=0.01$ with 50 projected GD iterations and a stepsize of 0.05. For the seismic deconvolution and MRI reconstruction problem, $\epsilon=1$, with 20 projected GD iterations and a stepsize of 0.1. AT is significantly slower than other methods. Notice that the jittering in SGD and SPGD does introduce very minimal time overhead compared to standard MSE training, even when accounting for the additional noise sampling steps.